How do you graph 7x+y=-2????
Please help!!! (this is not a bad question, it is asking for help)!!!
Alpha

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

How do you graph 7x+y=-2?
--------
The graph is a straight, so 2 points are needed.
7x+y=-2
y = -7x - 2
Pick 2 values for x and find y.
(0,-2) is an easy one.
Plot the 2 points.
Draw a line thru them.

Answer by ewatrrr(24785)


Hi, understand
Note: the standard slope-intercept form for an equation of a line: y = mx + b
where m is the slope and b the y-intercept.
7x+y = -2 |solving for y so as to put into the slope-intercept form
y = -7x -2 |slope m = -(7/1), y-intercept is Pt(0,-2}
Note: as y increases/decreases by 7, x value changes by 1 in the other direction
Pt(-1,5) and Pt(1,-9) on the line for example